# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ IV-B Resolvability Analysis
$\begin{split}&\mathbb{P}(\mathcal{E}^{c})=\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})W^{n}_{2}(\bm{Z}|\bm{X})}\bigg{(}\log\frac{W_{2}^{n}(\bm{Z}|\bm{X})}{\bar{Q}^{(n)}_{1}(\bm{Z})}\geq\gamma_{1}\bigg{)}\\ =&\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})}\left[\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg{(}\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{\bar{f}_{1}^{(n)}(\bm{Z})}\geq\gamma_{1}\bigg{)}\right].\end{split}$ (70)
Note that in (70), the distribution $\bar{Q}^{(n)}_{1}$ is the induced distribution of the channel $W_{2}^{n}$ by $P_{n}^{(n)}$ but not $\bar{P}_{n}^{(n)}$, which is a isotropic Gaussian distribution. Similarly as (42), we assume a codeword $\bm{x}_{0}=[\sqrt{\Upsilon},\cdots,\sqrt{\Upsilon}]$, where $\Upsilon\in[\mu^{2}\Psi,\Psi]$. The information density $\log\frac{W_{2}^{n}(\bm{z}|{\bm{x}_{0}})}{\bar{f}_{1}^{(n)}(\bm{z})}$ is expressed as
$\begin{split}&\frac{n}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})\\ +&\frac{\log e}{2}\sum_{i=1}^{n}\left[\frac{z_{i}^{2}}{\sigma^{2}+\mu\Psi(n)}-\frac{(z_{i}-\sqrt{\Upsilon})^{2}}{\sigma^{2}}\right]\\ =&\frac{n}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})-\sum_{i=1}^{n}\frac{1}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}\\ \times&\left[\mu\Psi(n)z_{i}^{2}-2(\sigma^{2}+\mu\Psi(n))\sqrt{\Upsilon}z_{i}+(\sigma^{2}+\mu\Psi(n))\Upsilon\right]\end{split}$ (71)
@@
$\bar{H}^{\Upsilon}_{n}=$ $\frac{n}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})$
@@
where we have used $\xi_{i}\sim\mathcal{N}(0,1)$ for $i=1,\cdots,n$. We rewrite (IV-B) as $n\ \mathsf{C}(\mu,\Psi(n),\sigma^{2})+\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})$, where $\mathsf{C}(\mu,\Psi(n),\sigma^{2})=\frac{1}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})$ and
$\begin{split}&S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})\\ =&\frac{\log e}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}\left(\sigma^{2}\Upsilon-\mu\Psi(n)\sigma^{2}\xi_{i}^{2}+2\sigma^{3}\sqrt{\Upsilon}\xi_{i}\right).\end{split}$ (73)
It is obvious that $S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2}),i=1,\cdots,n$ are i.i.d random variables and for each $i$ and any $\Upsilon\in[\mu^{2}\Psi,\Psi]$
$\begin{split}\mathbb{E}\left[S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})\right]=&\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}(\Upsilon-\mu\Psi(n))\\ =&\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2}).\end{split}$ (74)
Furthermore, we have
$\begin{split}\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})&\leq\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\\ &\,\ \ \ \ \ \ \ \ \ \,\forall\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)].\end{split}$ (75)
Now, let $\gamma_{1}$ be as in (66) with $\rho>0$. The quantity $\gamma_{1}$ does not depend on $\Upsilon$. By Chernoff bounding technique, we have
$\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x}_{0})}\bigg{(}\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}_{0}})}{\bar{f}_{1}^{(n)}(\bm{Z})}\geq\gamma_{1}\bigg{)}$
$\overset{(a)}{\leq}$ $\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x}_{0})}\left[\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}_{0}})}{\bar{f}_{1}^{(n)}(\bm{Z})}\geq n(1+\rho)\right.$
$\times$ $\left.[\mathsf{C}(\mu,\Psi(n),\sigma^{2})+\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})]\right]$
$\leq$ $ae^{-bn}\,\ \ \,\text{for some}\,\ \,a,b>0\,\,\text{depend on}\,\ \,\rho,$ (76)
where the step (a) follows from (75). Note that the above inequality holds for any $\Upsilon\in[\mu^{2}\Psi,\Psi]$. From sphere symmetry, for any $\bm{x}$ in the support of $\bar{P}_{n}^{(n)}$, the following inequality holds.
$\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg{(}\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{\bar{f}_{1}^{(n)}(\bm{Z})}\geq\gamma_{1}\bigg{)}\leq ae^{-bn}.$ (77)
@@
$\begin{split}\mathbb{P}(\mathcal{E}^{c})=&\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})W^{n}_{2}(\bm{Z}|\bm{X})}\bigg{(}\log\frac{W_{2}^{n}(\bm{Z}|\bm{X})}{\bar{Q}^{(n)}_{1}(\bm{Z})}\geq\gamma_{1}\bigg{)}\\ \leq&ae^{-bn}.\end{split}$ (78)
@@
$\mathbb{E}(\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{(n)}_{1}))\leq\log\big{(}1+\frac{2^{\gamma_{1}+1}}{MK}\big{)}+An\cdot ae^{-bn}.$ (79)
@@
$\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{X})\bar{P}_{n}^{(n)}(\bm{X})}\left(\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\hat{Q}^{(n)}_{1}(\bm{Z})}\right)\right)$
$=$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\hat{Q}^{(n)}_{1}(\bm{Z})}\right)\bigg{|}\mathcal{E}\right]\mathbb{P}(\mathcal{E})+\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\hat{Q}^{(n)}_{1}(\bm{Z})}\right)\bigg{|}\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
$=$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\hat{Q}^{(n)}_{1}(\bm{z})}\cdot\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{1}(\bm{Z})}\right)\bigg{|}\mathcal{E}\right]\mathbb{P}(\mathcal{E})$
$+$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\hat{Q}^{(n)}_{1}(\bm{z})}\cdot\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{1}(\bm{Z})}\right)\bigg{|}\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
$\overset{(a)}{\leq}$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{1}{(1-\iota^{\prime})}\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{1}(\bm{Z})}\right)\bigg{|}\mathcal{E}\right]\mathbb{P}(\mathcal{E})$
$+$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{1}{(1-\iota^{\prime})}\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{1}(\bm{Z})}\right)\bigg{|}\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
@@
In (IV-B), the inequality (a) follows by Lemma 3, the inequality (b) follows from the definition of $\mathcal{E}$, $\frac{1}{1-\iota^{\prime}}\leq 2$ for sufficiently $n$ and Lemma 7, (c) holds because of Lemma 8 and $\log\bigg{(}1+\frac{2}{MK}\cdot e^{\frac{n}{2\mu}}\big{(}1+\frac{\mu\Psi(n)}{\sigma^{2}}\big{)}^{\frac{n}{2}}\bigg{)}\leq An$ for some constant $A$. ∎
At this end, we have the following upper bound of $\mathbb{E}\left[\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{n}_{0})\right]$.
@@
$\begin{split}\mathbb{E}&\left[\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{n}_{0})\right]\leq\log\big{(}1+\frac{2^{\gamma_{1}+1}}{MK}\big{)}+An\cdot ae^{-bn}\\ +&\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})+(1+\iota^{\prime})\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})\end{split}$ (81)

### v5

@@ III-B Resolvability Analysis
$\begin{split}&\mathbb{P}(\mathcal{E}^{c})=\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})W^{n}_{2}(\bm{Z}|\bm{X})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|\bm{X})}{{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}>\gamma_{1}\bigg)\\ =&\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})}\left[\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)\right].\end{split}$ (76)
Note that in (76), the distribution ${Q}^{(n)}_{\bm{Z},1}$ is the induced distribution of the channel $W_{2}^{n}$ by $P_{n}^{(n)}$ but not $\bar{P}_{n}^{(n)}$, which is an isotropic Gaussian distribution. The inner term in (76)
$\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)$ (77)
is considered for each $\bm{x}_{sw}$ in the support set of $\bar{P}_{n}^{(n)}$. Similarly as the proof of Lemma 4, we first consider the probability of error in (77) for $\bm{x}_{sw}$ on the $n$ dimensional sphere with radius $\sqrt{n\Upsilon}$. The sphere symmetry allows us to assume $\bm{x}_{0}=[\sqrt{\Upsilon},\cdots,\sqrt{\Upsilon}]$. we assume a codeword $\bm{x}_{0}=[\sqrt{\Upsilon},\cdots,\sqrt{\Upsilon}]$, where $\Upsilon\in[\mu^{2}\Psi,\Psi]$. The information density $\log\frac{W_{2}^{n}(\bm{z}|{\bm{x}_{0}})}{h_{1}^{(n)}(\bm{z})}$ is expressed as
$\begin{split}&\frac{n}{2}\log\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)\\ +&\frac{\log e}{2}\sum_{i=1}^{n}\left[\frac{z_{i}^{2}}{\sigma^{2}+\mu\Psi(n)}-\frac{(z_{i}-\sqrt{\Upsilon})^{2}}{\sigma^{2}}\right]\\ =&\frac{n}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})-\sum_{i=1}^{n}\frac{1}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}\\ \times&\left[\mu\Psi(n)z_{i}^{2}-2(\sigma^{2}+\mu\Psi(n))\sqrt{\Upsilon}z_{i}+(\sigma^{2}+\mu\Psi(n))\Upsilon\right]\end{split}$ (78)
@@
$\bar{H}^{\Upsilon}_{n}=$ $\frac{n}{2}\log\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)$
@@
where we have used $\xi_{i}\sim\mathcal{N}(0,1)$ for $i=1,\cdots,n$. We rewrite (III-B) as $n\ \mathsf{C}(\mu,\Psi(n),\sigma^{2})-\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})$, where $\mathsf{C}(\mu,\Psi(n),\sigma^{2})=\frac{1}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})$ and
$\begin{split}&S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})\\ =&\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}\left(\mu\Psi(n)\xi_{i}^{2}-\Upsilon-2\sigma\sqrt{\Upsilon}\xi_{i}\right)\\ =&\frac{\mu\Psi(n)\log e}{2(\sigma^{2}+\mu\Psi(n))}\left[(\xi_{i}-\frac{\sigma\sqrt{\Upsilon}}{\mu\Psi(n)})^{2}-\frac{(\sigma^{2}+\mu\Psi(n))\Upsilon}{\mu^{2}\Psi^{2}(n)}\right]\\ =&\frac{\mu\Psi(n)\log e}{2(\sigma^{2}+\mu\Psi(n))}\theta_{i}^{2}-\frac{\Upsilon\log e}{2\mu\Psi(n)},\end{split}$ (80)
with $\theta_{i}=\xi_{i}-\frac{\sigma\sqrt{\Upsilon}}{\mu\Psi(n)}$. It is obvious that $S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2}),i=1,\cdots,n$ are i.i.d random variables. For each $i$ and any $\Upsilon\in[\mu^{2}\Psi,\Psi]$
$\begin{split}\mathbb{E}\left[S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})\right]=&\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}(\mu\Psi(n)-\Upsilon)\\ =&\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2}).\end{split}$ (81)
Furthermore, we have $\forall\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$,
$\begin{split}\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\leq&\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})\\ \leq&\hat{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n),\sigma^{2}),\end{split}$ (82)
and
$\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})=\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}(\mu-1)\Psi(n)<0.$ (83)
Now, let $\gamma_{1}$ be as in (72) with $\rho>0$, which does not depend on $\Upsilon$. The inequality (77) can be rewritten as (III-B), where the step (a) follows from (80), and step (b) follows from the tail probability bound of chi square distribution in (Theorem 4 in [32]), (82) and (83). The step (b) can also be derived by applying Chernoff bounding technique for sub-exponential random variables [33].
$\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x}_{0})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}_{0}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)$
$=$ $\mathbb{P}\left[n\ \mathsf{C}(\mu,\Psi(n),\sigma^{2})-\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})>n\mathsf{C}(\mu,\Psi(n),\sigma^{2})-n(1+\rho)\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\right]$
$=$ $\mathbb{P}\left[\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})<n(1+\rho)\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\right]$
$=$ $\mathbb{P}\left[\sum_{i=1}^{n}\left(S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})-\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})\right)<n\left(\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})-\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})\right)+n\rho\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\right]$
$\overset{(a)}{=}$ $\mathbb{P}\left[\frac{\mu\Psi(n)\log e}{2(\sigma^{2}+\mu\Psi(n))}\sum_{i=1}^{n}\left(\theta_{i}^{2}-E(\theta_{i}^{2})\right)<n\left(\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})-\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})\right)+n\rho\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\right]$
$\overset{(b)}{<}$ $a(\Upsilon,\rho)e^{-b(\Upsilon,\rho)n}\,\ \ \,\text{for some}\,\ \,a(\Upsilon,\rho),b(\Upsilon,\rho)>0$ (84)
Note that the above inequality (III-B) holds for any $\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$. Let $a=\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}\{a(\Upsilon,\rho)\}$ and $b=\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\min}\{b(\Upsilon,\rho)\}$, and both $a$ and $b$ are positive from the fact that $a(\Upsilon,\rho)$ and $b(\Upsilon,\rho)$ are continuous functions of $\Upsilon$. From sphere symmetry, for any $\bm{x}$ in the support of $\bar{P}_{n}^{(n)}$, the following inequality holds.
$\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)<ae^{-bn},$ (85)
@@
$\begin{split}\mathbb{P}(\mathcal{E}^{c})=&\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})W^{n}_{2}(\bm{Z}|\bm{X})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|\bm{X})}{{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}>\gamma_{1}\bigg)<ae^{-bn}.\end{split}$ (86)
Remark 9.
The strategy of the proof of Lemma 8 is similar as Lemma 4. We first consider the codewords with the same $l^{2}$ norm $\sqrt{n\Upsilon}$ (on the sphere with the same radius or with equal power constraint), separately. For codewords on each sphere, we can utilize the sphere symmetry to simplify the analysis. Again, the parameter $\mu$ makes the interval under control so that an universal threshold $\gamma_{1}$ can be obtained. The inner probability $\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)$ for each $\bm{x}$ is a tail probability and can be bounded by concentration inequality from Chernoff technique. As the bounds of the tail probability depends only on the radius of the codewords and are all exponentially decreasing with $n$, we can get the maximum of the bounds for all $\bm{x}$, which is again exponential decreasing with $n$.
@@
$\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|\bar{Q}^{(n)}_{\bm{Z},1})\right]\leq\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)+An\cdot ae^{-bn}.$ (87)
@@
The details can be found in (III-B), where the first step follows from (70). The inequality (a) follows by Lemma 3, the inequality (b) follows from the definition of $\mathcal{E}$, $\frac{1}{1-\iota^{\prime}}\leq 2$ for sufficiently large $n$ and Lemma 7, (c) holds because of Lemma 8 and $\log\bigg(1+\frac{2}{MK}\cdot e^{\frac{n}{2\mu}}\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)^{\frac{n}{2}}\bigg)\leq An$ for some constant $A>0$. ∎
$\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|\bar{Q}^{(n)}_{\bm{Z},1})\right$]
$\leq$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{X})\bar{P}_{n}^{(n)}(\bm{X})}\left(\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\right)$
$=$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}\right]\mathbb{P}(\mathcal{E})+\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
$=$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}\cdot\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MKQ^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}\right]\mathbb{P}(\mathcal{E})$
$+$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}\cdot\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MKQ^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
$\overset{(a)}{\leq}$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{1}{(1-\iota^{\prime})}\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MKQ^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}\right]\mathbb{P}(\mathcal{E})$
$+$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{1}{(1-\iota^{\prime})}\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MKQ^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
@@
At this end, we have the following upper bound of $\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\right$] by upper bounding the second and third terms of (69).
@@
$\begin{split}\mathbb{E}&\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\right]\leq\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)+An\cdot ae^{-bn}\\ +&\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})\end{split}$ (89)